The amplitude of a driven harmonic oscillator reaches a value of 28.6 F0/m at a resonant frequency of 382Hz. What is the Q value of this system?


1
Expert's answer
2021-03-31T09:39:03-0400
"28.6=\frac{Q}{(\omega)^2}\\28.6=\frac{Q}{(2\pi (382))^2}\\Q=1.65\cdot10^8"